The following information is available for Hildebrand Corporation: Dividends paid to common stockholders Dividends paid to preferred stockholders Net income Weighted average common shares outstanding $45,000 20,000 295,000 100,000 Compute the earnings per share of common stock. (Round answer to 2 decimal places, e.g. 10.50,) Earnings per share Qu
Answer #1

EPS = earnings available to common stock holders/ weighted average number of common stock outstanding

earnings available to common stock holders =net income - preferred dividends = 295,000 - 20,000 = 275,000

EPS = 275,000/100,000 = 2.75 per share